Question**
An angle measure is 8 times the size of its supplement. Find the measure of the angle.

Answer by richwmiller(17219)

a+s=180
a=8s
a/8=s
a+a/8=180
8a+a=8*180
9a=8*180
a=8*20=160 =angle
s=20